OpenStudy (anonymous):

Laura is baking a cake. The recipe says that she has to mix 64 grams of chocolate powder to the flour. Laura knows that 1 cup of that particular chocolate powder has a mass of 128 grams. She added two over three of a cup of chocolate powder to the flour. Should Laura add more chocolate powder to make the exact recipe, or did she go over and by what amount? She went over by one over three of a cup She needs to add one over three of a cup She needs to add one over six of a cup She went over by one over six of a cup

OpenStudy (dn_20):

So x= 64, or the required amount of gram to the flour y= # of powder she add She add 2/3 of 128 in, so: 128y = 128 2/3 = 85.3 g 64 is 1/2 of 128, so: 2/3 -1/2 = 4/6 - 3/6 = 1/6 She went over by one over six of a cup

OpenStudy (anonymous):

Thank You so Much

OpenStudy (dn_20):

:))